# Quantifying individual and team performance in cricket

### v1

@@ II.2 Long-term performance from early career observations
Figure 3: Leadership affects player performance. (a) Boxplot showing the average runs scored and average wickets taken by captain and non-captain players. The median trend for batsmen and bowlers is opposite with captain-batsmen performing better than their non-captain counterparts. PDFs of performance for captains who are (b) batsman and (c) bowlers during various phases of their career. The dotted black lines denote the median of the distribution. The performance of captain-batsmen improves during captaincy, while it declines for captain-bowlers.
II.3 Effect of drop and re-entry
In the team-oriented sport of cricket, fluctuations in team composition frequently arise due to injuries and variations in player performance. While injuries often occur unpredictably, a decline in performance typically manifests more gradually and may not be immediately apparent. A fitting inquiry in this context is whether it is possible to discern patterns in a player’s performance preceding their expulsion from the team. Additionally, it is important to determine whether a player’s performance improves post-reinstatement to the team.
@@ II.3 Effect of drop and re-entry
Figures (2) (c) and (d) show the average performance trajectories of batsmen and bowlers, respectively, both prior to their removal from the team and subsequent to their rejoining. We consider all players who experienced a temporary exclusion from the team for a minimum of three matches before making a return. We observe a consistent downtrend in performance during the pre-removal phase, with average runs and wickets demonstrating a monotonic decline, amounting to an approximate 19% reduction for both batsmen and bowlers. Notably, the lowest performance levels are recorded in the match just before the player’s exclusion. By contrast, players tend to exhibit a strong comeback performance after reinstatement to the team. In particular, the initial performance post-return exhibits a substantial elevation, with batsmen registering an approximate 36% improvement and bowlers showing a 30% enhancement compared to their final pre-removal performance. This enhanced performance appears to be stable in subsequent matches, consistently displaying higher performance compared to the pre-removal, suggesting that short temporal exclusion may improve performance in the longer run.
II.4 Role of leadership
In team sports, the composition of an effective team encompasses skilled players under proficient leadership. Case studies across various domains, including sports, science, and business, have illustrated scenarios, where moderately weak teams achieve success under adverse conditions, guided by strategic leadership Guimerà et al. (2005); Zhang et al. (2022); Hancock et al. (2023). Here we explore the influence of captaincy on a player’s performance. Specifically, we investigate whether the added responsibility of leadership alters player performance and if such an impact persists throughout their careers.
We consider captains who have led their team in a minimum of 50 matches (see Methods). Within our dataset, we identify 71 captains, comprising 69% batsmen and 31% bowlers, indicating a higher propensity for batsmen to be chosen as captains – nearly twice as often as bowlers. In fact, many of the most successful captains including Allan Border, Graeme Smith, M S Dhoni, Ricky Ponting, and Sourav Ganguly have been batsmen. However, only a handful of captains such as Imran Khan, Shaun Pollock, and Wasim Akram have been bowlers. Despite these differences, captain-bowlers remain in this position for longer than the batsmen on average (see SM4).
Figure (3) (a) presents a comparison of the career average performances between captains and non-captains (referred to as players). A detailed analysis reveals a notable variation in performance levels between captain-batsmen and captain-bowlers, compared to their non-captain counterparts. Quantitatively, the average runs scored by captain-batsmen ($32$) exceed those of players ($25$) by approximately 20%. Conversely, captain-bowlers secure, on average, 21% fewer wickets than non-captain-bowlers ($0.9$ versus $1.1$). We have considered only the matches where players get a chance to perform. This disparity suggests differing pathways to captaincy, where batsmen may need to consistently outperform peers, while bowlers’ captaincy seems less dependent on individual performance.
To further elucidate this phenomenon, we divide the careers of captains into three phases: pre-captaincy, during captaincy, and post-captaincy, for both batsmen (fig (3) (b)) and bowlers (fig (3) (c)). We observe that batsmen typically enhance their performance upon assuming captaincy, whereas bowlers exhibit a decline. Closer examination of fig (3) (c) reveals a bimodal distribution during the captaincy phase for bowlers, indicating a dichotomy in skill sets among captain-bowlers. Post-captaincy, the trajectories diverge significantly for batsmen and bowlers. For batsmen, average performance markedly decreases post-captaincy, even falling below their pre-captaincy levels. In contrast, bowlers tend to recover and often improve upon their captaincy phase performance. These observations suggest a close correlation between individual performance and captaincy tenure for batsmen, where they often lead through exemplary performance. Conversely, bowlers may experience captaincy as a burden, reflected in their individual statistics. This could account for the observed differences in performance between captains and non-captains among batsmen and bowlers in fig. (3) (a).
II.5 Contribution of specialists
@@
Figure 4: Contribution of specialists. (a) Runs and strike rates of opener and non-opener batsmen. The solid black lines denote the mean of the distributions. Openers score more runs at a slower pace, while non-openers play aggressively at the cost of lower contribution to team scores. (b) Fractional contribution $f_{c}$ of players to the team scores for batsman, bowler, and all-rounder. All-rounders contribute significantly higher to the team. (c) Fractional dismissals $d_{i}$ per match by wicketkeeper and other fielders over the entire career. Wicketkeepers exhibit a very high tendency to contribute to team dismissals as compared to the fielders.
In team sports, the strategic composition of a team is crucial for its effective functioning. The concept of utilizing specialists – individuals who perform highly specific roles – extends beyond sports into various domains of human activity, including business organizations, scientific research, software development and even hunter-gatherer societies. In this study, we examine three distinct specialist roles within cricket, encompassing different facets of the game: namely, opening batsmen, all-rounders (contributing with both bat and ball), and wicket-keepers.
Figure (4) (a) elucidates the variations in batting patterns between opening batsmen and their non-opening counterparts. Batsmen are categorized as openers if they occupy the first or second position in the batting lineup, while positions 3 to 6 are considered non-openers (see Methods). We focus on the pace at which they accumulate these runs. The run-scoring rate is quantified as the number of runs per 100 balls faced, commonly referred to as the strike-rate in cricket terminology. This metric is frequently used to assess a player’s defensive or aggressive playing style. Virender Sehwag (average strike-rate=94) and Brendon McCullum (average strike-rate=97) are prime examples of explosive openers who scored runs at a brisk pace in their career. However, our analysis reveals that openers average a strike rate of around 63 runs per 100 balls, whereas non-openers exhibit a higher average of 68 runs per 100 balls. This suggests that non-openers score approximately 9% more rapidly than their opening counterparts. However, the average score for openers stands at approximately 31 runs, in contrast to 26 runs for non-openers, indicating a reduction of about 17% for batsmen coming later in the batting order. Collectively, these findings illustrate a strategic balance between defensive and attacking approaches contingent on a player’s batting position. Openers tend to adopt a more conservative approach, possibly due to the initial uncertainty of the pitch conditions, while subsequent batsmen often adopt a more aggressive strategy, building upon the foundational efforts of the openers. Further analysis reveals that this disparity may be partially attributed to the differing number of balls faced by players based on their batting order (see SM5).
@@
All-rounders are players who make significant contributions in both aspects of the game – batting and bowling (see Methods). In contrast, specialized batsmen and bowlers are characterized by their significant contribution in only one of these areas. To quantitatively assess the impact of each player type, we introduce the fractional contribution $f_{c}$ as,
, so that a player’s maximum possible contribution ($f_{c}$) is normalized within the range of 0 to 1. It is important to note that, given this framework, specialized batsmen and bowlers have an upper limit of $f_{c}=0.5$ in best-case scenario. Figure (4) (b) presents the distribution of the contributions of all players across all matches. Among these, batsmen demonstrate the lowest mean contribution ($f_{c}\approx 0.06$), with bowlers exhibiting a slightly higher average contribution of $f_{c}\approx 0.1$. Notably, all-rounders surpass both groups with an average contribution of $f_{c}\approx 0.11$. Furthermore, there are instances where all-rounders singularly account for more than 50% of the team’s performance, evidenced by instances of $f_{c}>0.5$. We also note that non-specialists contribute significantly less than the specialists in both departments (see SM6).
In summary, while bowlers generally contribute more towards the team’s collective efforts than batsmen, all-rounders emerge as pivotal players, often driving the team’s success with contributions that exceed 50% of the total team scores. This finding underscores the vital role of all-rounders in cricket. Future research could delve into a more detailed examination of the all-rounder’s role and their strategic impact on the dynamics of the game.
We now concentrate on the impact of fielding (see Methods) in cricket. Although most players are predominantly recognized for their batting or bowling abilities, fielding is an integral aspect of a team’s overall success. Indeed, a commonly reiterated phrase in cricket states, ‘catches win matches’, emphasizing the significance of player dismissals through effective fielding. Among various fielding roles, the wicket-keeper is considered to hold a pivotal position (see Methods). Typically, a batsman can be dismissed in three primary ways: (a) if the bowler successfully hits the stumps or the batsman obstructs the stump with their leg, (b) if a fielder catches the ball hit by the batsman before it touches the ground, or (c) if the stump is dislodged by the fielders before the completion of a run between the wickets  [^1] There are other forms of dismissals such as ‘retired hurt’, ‘obstructing fielding’, etc. but we do not consider them since the frequency of such dismissals is low.. The first method relies exclusively on the bowler’s skill, whereas the latter two require strategic collaboration between the bowler and fielders. Our interest lies in the last two dismissal types. Specifically, we define $d_{i}$ as the number of dismissals executed by a fielder of type (b) and (c) per match. Figure (4) (c) presents a boxplot depicting the distribution of dismissal fractions attributed to wicket-keepers and fielders. Our observations reveal that wicket-keepers play a vital role in dismissals, accounting for approximately 0.7 dismissals per match. In contrast, fielders contribute to a relatively minor number of dismissals, averaging around 0.3. The strategic positioning of wicket-keepers behind the wickets significantly facilitates these outcomes with well-known captains such as M.S. Dhoni (441 dismissals in 340 matches) winning multiple accolades for India and Kumar Sangakkara (477 dismissals in 391 matches) guiding Sri-Lanka to the finals of the 2011 World Cup.
@@
Figure 5: Patterns in team success (a) Occurrences of hot streaks (wins denoted by red) or cold streaks (losses denoted by blue) of a given length $L_{streak}$ compared to a null model. $N(l\geq L_{streak}$ denotes the number of times when the team had the same result for $l$ or more matches, averaged over teams. We observe that teams demonstrate both hot and cold streaks. The fluctuations in the numbers calculated as the standard error are smaller than the symbols. (b) Differences between effective team size $S_{eff}$ for winning and losing teams. Winning teams have a larger $S_{eff}$ on average indicating that a larger number of players contribute significantly in winning teams than in losing teams.
In summary, team success depends on the successful coordination of different types of – often specialized – contributions. Openers, by adopting a cautious approach assessing pitch conditions and scoring at a lower rate, lay a strong foundation for the team. All-rounders enhance the team’s score by applying their skills across both facets of the game. Lastly, wicket-keepers, by contributing to a substantial number of total dismissals, enhance the bowlers’ efforts and, consequently, the overall team performance.
II.6 Patterns of team success
So far we have predominantly focused on individual contributions to team success in cricket. We now broaden our perspective to analyze team success, setting aside specific individualistic factors that contribute to victory. Given that each cricket match culminates in a definitive outcome (see Methods for more information on what it means to win), our interest lies in discerning patterns of wins and losses for teams.
@@ II.6 Patterns of team success
Figure (5) (a) outlines the occurrence of hot streaks within teams. Similar to the bursts of exceptional performance observed in individual players (as discussed in sec.(II.1)), we assess the propensity of teams to consecutively win (or lose) a series of matches. A familiar example is Australia’s 21 match winning streak in 2003 which culminated in them winning the World Cup. We quantify this tendency by defining the ratio $\frac{N_{data}}{N_{rand}}$, where $N_{data}$ represents the frequency of a team winning (or losing) $l$ or more consecutive matches. In contrast, $N_{rand}$ is the analogous frequency within a randomized null model (see Methods). Thus, the ratio $\frac{N_{data}}{N_{rand}}$ indicates the likelihood of a team experiencing a streak of at least $l$ consecutive wins or losses compared to a random outcome. Figure (5) (a) reveals that teams are more prone to winning or losing matches in sequence than would be expected by chance. Notably, the probability of winning 7 or more consecutive matches is nine times higher than chance, while the likelihood of losing 7 or more consecutive matches is about three times the random expectation. This observation suggests that teams exhibit hot streaks in their winning performances. Intriguingly, we also observe cold streaks, where teams undergo multiple successive losses.
We then characterise the relation between collective contributions and match outcomes. Specifically, we employ the concept of effective team size, a well-established metric in the team analysis literature Klug and Bagrow (2016); Delice et al. (2019). Effective team size quantifies the essential number of players in a team, effectively measuring the redundancy in team composition. Mathematically, we denote the contribution of the player to the team score as $f_{c}$, such that $\sum_{c}f_{c}=1$, as per Eq. (1). The effective team size $S_{eff}$ is defined in such a way that it equals the actual team size if contributions are evenly distributed among all players and reduces to 1 if a single player is solely responsible for the entire performance. A common approach to calculate $S_{eff}$ is to use the formula $S_{eff}=2^{H}$, where $H=-\sum_{c}f_{c}\log_{2}f_{c}$. Statistically, $S_{eff}$ represents the entropy of the distribution of contributions $f_{i}$ from team players.
As illustrated in Figure (5) (b) via a boxplot, the distributions of $S_{eff}$ for winning and losing teams differ notably. The median $S_{eff}$ for winning teams ($\approx 7.3$) is approximately 5% greater than that for losing teams ($\approx 6.9$). This indicates that successful teams require a broader array of individual contributions. Indeed in the 2015 World Cup, Australia, the winner, had $S_{eff}=6.77$ in the finals, as compared to New-Zealand who had $S_{eff}=5.6$. Surprisingly, in the controversial final of the 2019 World Cup, despite being the losing side, New-Zealand had $S_{eff}=8.1$, while England had $S_{eff}=7.7$. Moreover the data also reveal a redundancy of 3 to 4 players in most teams. However, our supplementary analysis also reveals that the top performers from the winning team contribute more to their team than their top performing counterparts from the losing team (see SM7).
In conclusion, our analysis reveals that teams tend to win or lose matches in clusters. However, for a team to secure victory, it is imperative that multiple individuals make substantial contributions. A lack of such collective effort often results in teams losing matches.
III Discussion
@@
In the last few years, the advent of detailed sports analytics has revolutionised our understanding of human behaviour in sports. It offers insights into physical performance, cognitive processes, and team dynamics, moving beyond traditional training methods. This shift towards a data-driven approach in sports reflects a broader societal trend in optimizing human potential, combining historical practices with modern scientific methodologies.
@@
In this study, we quantified the individual and team performance in cricket, examining various aspects of the game. Due to the absence of quantifiable performance metrics such as the Elo rating system in chess, we validated the use of runs scored and wickets taken as indicators of batting and bowling performance, respectively. Our analysis revealed a significant increase in average runs scored over time. To negate this inflation in run-scoring and facilitate fair comparisons, we implemented a normalization of all performances involving runs. This rescaled data corroborated the “random impact rule” observed in other domains, indicating that a player’s peak performance can occur at any point during their career. However, this best performance often co-occurs with other comparable exemplary performances, pointing to the existence of hot streaks in individual cricket players. Our exploration into the predictors of individual performance unveiled a strong correlation between early career achievements and overall career trajectory. Additionally, we confirmed that players are often excluded from teams due to declining performances, yet they typically exhibit a sustained enhancement in performance following their comeback.
We investigated the relationship between player performance and captaincy. Our data suggested that batsmen who ascend to captaincy roles demonstrate higher performance levels, both overall and during their tenure as captain, in contrast to non-captain-batsmen. This pattern was not mirrored among bowlers, highlighting the differential impact of captaincy on distinct player types. Focusing on the role of specialists, our findings indicated that opening batsmen typically score more runs but at a slower rate, whereas subsequent batsmen tend to adopt more aggressive strategies. All-rounders were observed to contribute significantly more to team performance compared to batsmen. In the realm of fielding, wicketkeepers emerged as pivotal in effecting dismissals, thereby bolstering the impact of bowlers on team success. Additionally, our analysis identified both hot and cold streaks within teams, underscoring that victories and defeats tend to occur in clusters. Winning teams were characterized by more evenly distributed player performances, as indicated by higher effective team sizes.

### v2

@@ III.2 Long-term performance from early career observations
Figure 3: Leadership affects player performance. (a) Boxplot showing the average runs scored and average wickets taken by captain and non-captain players. The average trend for batsmen and bowlers is the opposite with captain-batsmen performing better than their non-captain counterparts. PDFs of performance for captains who are (b) batsman and (c) bowlers during various phases of their career. The dotted black lines denote the median of the distribution. The performance of captain-batsmen improves during captaincy, while it declines for captain-bowlers.
III.3 Effect of drop and re-entry
Persistence (or lack thereof) of team composition can have a substantial effect on individual as well as team performance. Methods: Figures (2) (c) and (d) show the average performance trajectories of batsmen and bowlers, respectively, both prior to their removal from the team and subsequent to their rejoining. We consider all players who experienced a temporary exclusion from the team for a minimum of three matches before making a return.
@@ III.3 Effect of drop and re-entry
Results: We observe a consistent downtrend in performance during the pre-removal phase, with average runs and wickets demonstrating a monotonic decline, amounting to an approximate 19% reduction for both batsmen and bowlers compared to 5 matches before the exclusion. Notably, the lowest performance levels are recorded in the match just before the player’s exclusion. By contrast, players tend to exhibit a strong comeback performance after reinstatement to the team. In particular, the initial performance post-return exhibits a substantial elevation, with batsmen registering an approximate 36% improvement and bowlers showing a 30% enhancement compared to their final pre-removal performance. This enhanced performance appears to be stable in subsequent matches, consistently displaying higher performance compared to the pre-removal, suggesting that short temporal exclusion may improve performance in the longer run.
III.4 Role of leadership
Proficient leadership can enhance performance in teams of skilled individuals. Methods: We consider captains who have led their team in a minimum of 15 matches (Sec. (II.2). Within our dataset, we identify 172 captains, comprising 71% batsmen and 29% bowlers, indicating a higher propensity for batsmen to be chosen as captains – more than twice as often as bowlers. Despite these differences, captain-bowlers remain in this position for longer than the batsmen on average (see SM5). We first compare the career performance of captains with other players for both batsmen and bowlers. We use the Mann-Whitney $U$ test to determine if the distribution of captain performances is the same as that of the players’ performance distribution. To investigate the influence of leadership, we divide the career trajectories of captains into 3 phases, before, during, and after the captainship. We are interested if the performance qualitatively changes during the three phases.
@@ III.4 Role of leadership
Figure 4: Contribution of specialists. (a) Runs and strike rates of opener and non-opener batsmen. The solid black lines denote the mean of the distributions. Openers score more runs at a slower pace, while non-openers play aggressively at the cost of lower contribution to team scores. (b) Fractional contribution $f_{c}$ of players to the team scores for batsman, bowler, and all-rounder. All-rounders contribute significantly higher to the team. (c) Fractional dismissals $d_{i}$ per match by wicket-keeper and other fielders over the entire career. wicket-keepers exhibit a very high tendency to contribute to team dismissals as compared to the fielders.
Results: Figure (3) (a) presents a comparison of the career average performances between captains and non-captains (referred to as players). Quantitatively, the average runs scored by captain-batsmen ($31$) exceed those of players ($26$) by approximately 16%. Conversely, captain-bowlers secure, on average, 18% fewer wickets than non-captain-bowlers ($0.96$ versus $1.13$). We have considered only the matches where players get a chance to perform. For both batsmen and bowler, the Mann-Whitney $U$ test yields $p$-value $<0.001$. This disparity suggests differing pathways to captaincy, where batsmen may need to consistently outperform peers, while bowlers’ captaincy seems less dependent on individual performance.
For both batsmen (fig (3) (b)) and bowlers (fig (3) (c)). We observe that batsmen typically enhance their performance upon assuming captaincy, whereas bowlers exhibit a decline. A closer examination of fig (3) (c) reveals a bimodal distribution during the captaincy phase for bowlers, indicating a dichotomy in skill sets among captain-bowlers. Post-captaincy, the trajectories diverge significantly for batsmen and bowlers. For batsmen, average performance markedly decreases post-captaincy, even falling below their pre-captaincy levels. The decrease is even more substantial for bowlers.
III.5 Contribution of specialists
Specialists, providing focused contributions to group endeavours are often key for team success. Methods: Here we consider three types of specialists – openers, all-rounders, and wicket-keepers. Batsmen are categorised as openers if they occupy the first or second position in the batting lineup, while positions 3 to 6 are considered non-openers (Sec. (II.2)). We focus on the pace at which they accumulate these runs. The run-scoring rate is quantified as the number of runs per 100 balls faced, commonly referred to as the strike rate in cricket terminology. This metric is frequently used to assess a player’s defensive or aggressive playing style. We use the Mann-Whitney $U$ test to compare the distributions of runs (and strike rate) scored by openers and non-openers.
@@ III.5 Contribution of specialists
All-rounders are players who make significant contributions in both aspects of the game – batting and bowling (Sec. (II.2)). In contrast, specialised batsmen and bowlers are characterised by their significant contribution in only one of these areas. We quantify the fractional contribution for each type of player as shown in section (II.4.2). We use the K-S test to analyse the differences between the distributions of fractional contributions of batsmen and bowlers against all-rounders.
@@
Figure 5: Occurrences of hot streaks (wins denoted by red) or cold streaks (losses denoted by blue) of a given length $L_{streak}$ compared to a null model. $N(l\geq L_{streak}$ denotes the number of times when the team had the same result for $l$ or more matches, averaged over teams. We observe that teams demonstrate both hot and cold streaks. The fluctuations in the numbers calculated as the standard error are smaller than the symbols. (b) Differences between effective team size $S_{eff}$ for winning and losing teams. Winning teams have a larger $S_{eff}$ on average indicating that a larger number of players contribute significantly in winning teams than in losing teams.
Among various fielding roles, the wicket-keeper is considered to hold a pivotal position. Typically, a batsman can be dismissed in three primary ways: (a) if the bowler successfully hits the stumps or the batsman obstructs the stump with their leg, (b) if a fielder catches the ball hit by the batsman before it touches the ground, or (c) if the stump is dislodged by the fielders before the completion of a run between the wickets  [^1] There are other forms of dismissals such as ‘retired hurt’, ‘obstructing fielding’, etc. but we do not consider them since the frequency of such dismissals is low.. The first method relies exclusively on the bowler’s skill, whereas the latter two require strategic collaboration between the bowler and fielders. Our interest lies in the last two dismissal types. Specifically, we define $d_{i}$ as the number of dismissals executed by a fielder of type (b) and (c) per match. We use Welch’s t-test to compute the statistical significance of differences between the number of dismissals by the wicket-keepers and fielders respectively.
Results: Figure (4) (a) elucidates the variations in batting patterns between opening batsmen and their non-opening counterparts. Our analysis reveals that openers average a strike rate of around 63 runs per 100 balls, whereas non-openers exhibit a higher average of 69 runs per 100 balls. This suggests that non-openers score approximately 9% more rapidly than their opening counterparts. However, the average score for openers stands at approximately 31 runs, in contrast to 26 runs for non-openers, indicating a reduction of about 17% for batsmen coming later in the batting order. The Mann-Whitney $U$ test yields statistical significance level given by $p$-value $<0.001$.
Figure (4) (b) presents the distribution of the contributions of all players across all matches. Among these, batsmen demonstrate the lowest mean contribution ($f_{c}\approx 0.06$), with bowlers exhibiting a slightly higher average contribution of $f_{c}\approx 0.1$. Notably, all-rounders surpass both groups with an average contribution of $f_{c}\approx 0.11$. Furthermore, there are instances where all-rounders singularly account for more than 50% of the team’s performance, evidenced by instances of $f_{c}>0.5$. The K-S test reveals that the contributions from batsmen and bowlers are significantly different ($p$-value $<0.001$) from those of all-rounders. We also note that non-specialists contribute significantly less than the specialists in both departments (see SM8).
Figure (4) (c) presents a boxplot depicting the distribution of dismissal fractions attributed to wicket-keepers and fielders. Our observations reveal that wicket-keepers play a vital role in dismissals, accounting for approximately 0.7 dismissals per match. In contrast, fielders contribute to a relatively minor number of dismissals, averaging around 0.3. The Welch’s t-test gives a $p$-value $<0.001$ implying that wicket-keepers are consistently better than fielders when the opportunity arises.
III.6 Patterns of team success
@@
Methods: Similar to the bursts of exceptional performance observed in individual players (as discussed in Sec.(III.1)), we assess the propensity of teams to consecutively win (or lose) a series of matches. We quantify this tendency by defining the ratio $\frac{N_{data}}{N_{rand}}$, where $N_{data}$ represents the frequency of a team winning (or losing) $l$ or more consecutive matches. In contrast, $N_{rand}$ is the analogous frequency within a randomised null model (see Methods). Thus, the ratio $\frac{N_{data}}{N_{rand}}$ indicates the likelihood of a team experiencing a streak of at least $l$ consecutive wins or losses compared to a random outcome.
To gain more insight into the team composition, we employ the concept of effective size $S_{eff}$ as discussed in Methods. To compute the statistical differences between the $S_{eff}$ of winning and losing teams, we use the Mann-Whitney $U$ test. Additionally, we also compute the effect size $r=\frac{U}{n_{1}n_{2}}$, where $U$ is the test statistic and $n_{1}$ and $n_{2}$ are the sample sizes respectively.
Results: Figure (5) (a) reveals that teams are more prone to winning or losing matches in sequence than would be expected by chance. Notably, the probability of winning 7 or more consecutive matches is nine times higher than chance, while the likelihood of losing 7 or more consecutive matches is about three times the random expectation. This observation suggests that teams exhibit hot streaks in their winning performances. Intriguingly, we also observe cold streaks, where teams undergo multiple successive losses.
As illustrated in Figure (5) (b) via a boxplot, the distributions of $S_{eff}$ for winning and losing teams differ notably. The median $S_{eff}$ for winning teams ($\approx 6.9$) is approximately 4% greater than that for losing teams ($\approx 6.6$). This indicates that successful teams require a broader array of individual contributions. The Mann-Whitney $U$ test indicates that the differences are significant ($p$-value $<0.001$) and the effect size $r=0.56$ indicates a probability that the winning side is 56% more likely to have higher $S_{eff}$ than the losing side.
IV Discussion
@@
In the last few years, the advent of detailed sports analytics has revolutionised our understanding of human behaviour in sports Lewis (2004); Brown et al. (2017); Pappalardo and Cintia (2018). It offers insights into physical performance, cognitive processes, and team dynamics, moving beyond traditional training methods Mason and Foster (2007); Dawson (2023). This shift towards a data-driven approach in sports reflects a broader societal trend in optimising human potential, combining historical practices with modern scientific methodologies.
@@
In this study, we quantified the individual and team performance in ODI cricket, examining various aspects of the game. Due to the absence of quantifiable performance metrics such as the Elo rating system in chess, we validated the use of runs scored and wickets taken as indicators of batting and bowling performance, respectively. Our analysis revealed a significant increase in average runs scored over time. To negate this inflation in run-scoring and facilitate fair comparisons, we implemented a normalisation of all performances involving runs (see SM1).
This rescaled data corroborated the “random impact rule” observed in other domains, indicating that a player’s peak performance can occur at any point during their career Sinatra et al. (2016); Liu et al. (2018). However, this best performance often co-occurs with other comparable exemplary performances, pointing to the existence of hot streaks in individual cricket players. Across a range of domains such as science Liu et al. (2018, 2022), arts Williams et al. (2019); Janosov et al. (2020), and sports Ram et al. (2022); Chowdhary et al. (2023), individual careers display hot streaks, where exceptional performances tend to occur in bursts which are clustered in time. Our analysis showed that, while the best performance of a cricket player’s career may manifest at any point within their career, peak performances tend to occur in rapid sequence over a short period of time. This is further confirmed by the probability of performing well in consecutive matches growing exponentially as evidenced by the recurrence time distribution in SM3.
Our exploration into the predictors of individual performance unveiled a strong correlation between early career achievements and overall career trajectory. However noteworthy differences were observed between batsmen and bowlers with batsmen improving their performance from early careers more than the bowlers. Such disparities may be attributed to the distinct elements inherent in the roles of batsmen and bowlers. Batsmen, for instance, may leverage accumulated experience and refined skills to augment run-scoring, while bowlers may depend more on innovation and strategic creativity to increase their wicket tally. This distinction potentially accounts for the observed trend of a higher proportion of batsmen surpassing their initial career performance compared to bowlers. Alternatively, our observations can also be explained if some players are simply better than the rest. Thus, their performance in various stages of their careers will correlate with their full career performance. This effect is also known as the Q-model in the context of scientific careers Sinatra et al. (2016). Further analyses supporting this hypothesis are presented in SM4. Additionally, we confirmed that players are often excluded from teams due to declining performances, yet they typically exhibit a sustained enhancement in performance following their comeback.
We investigated the relationship between player performance and captaincy. Our data suggested that batsmen who ascend to captaincy roles demonstrate higher performance levels, both overall and during their tenure as captain, in contrast to non-captain-batsmen. This pattern was not mirrored among bowlers, highlighting the differential impact of captaincy on distinct player types. These observations suggest a close correlation between individual performance and captaincy tenure for batsmen, where they often lead through exemplary performance. Conversely, bowlers may experience captaincy as a burden, reflected in their individual statistics. This could account for the observed differences in performance between captains and non-captains among batsmen and bowlers. We tested that this phenomenon is not because the captain-bowlers bowl for more time in SM6. An alternative explanation might be that captains tend to bowl at the wrong time to relieve the team pressure. However, to test this hypothesis one would need fine-grained temporal information about the progression of a match, which is not present in our data.
Focusing on the role of specialists, our findings indicated that opening batsmen typically score more runs but at a slower rate, whereas subsequent batsmen tend to adopt more aggressive strategies. Collectively, these findings illustrate a strategic balance between defensive and attacking approaches contingent on a player’s batting position. Openers tend to adopt a more conservative approach, possibly due to the initial uncertainty of the pitch conditions, while subsequent batsmen often adopt a more aggressive strategy, building upon the foundational efforts of the openers. Further analysis reveals that this disparity may be partially attributed to the differing number of balls faced by players based on their batting order (see SM7).
All-rounders enhance the team’s score by applying their skills across both facets of the game. They were observed to contribute more to team performance compared to batsmen and bowlers. While bowlers generally contribute more towards the team’s collective efforts than batsmen, all-rounders emerge as pivotal players, often driving the team’s success with contributions that exceed 50% of the total team scores. In the realm of fielding, wicket-keepers emerged as pivotal in effecting dismissals, thereby bolstering the impact of bowlers on team success. In summary, team success depends on the successful coordination of different types of – often specialised – contributions. Although most players are predominantly recognised for their batting or bowling abilities, fielding is an integral aspect of a team’s overall success. Indeed, a commonly reiterated phrase in cricket states, ‘catches win matches’, emphasising the significance of player dismissals through effective fielding. Wicket-keepers, by contributing to a substantial number of total dismissals, enhance the bowlers’ efforts and, consequently, the overall team performance. These findings underscore the vital role of specialists in ODI cricket. Future research could delve into a more detailed examination of their role and their strategic impact on the dynamics of the game.
Additionally, our analysis reveals, similarly to individual performances, also teams tend to win or lose matches in clusters, thus identifying both hot and cold streaks during seasons. We observed that winning teams were characterised by more evenly distributed player performances, as indicated by higher effective team sizes. A lack of such collective effort often results in teams losing matches. Moreover the data reveals a redundancy of 3 to 4 players in most teams. However, our supplementary analysis also reveals that the top performers from the winning team contribute more to their team than their top performing counterparts from the losing team (see SM9).